The length of a rectangular garden is 4 yd greater than the width. The area of the garden is 96 〖yd〗^2.Find the length and the width.
Answer by mananth(16949)

width of garden = x yards
Length of garden = x+ 4 yards

Area = L*W
Area= 96 sq. yards
Area = x(x+ 4 )
x(x+ 4 )= 96

x^2 + 4 x = 96
x^2 + 4 x -96 = 0
x^2+ 12 x -8 x - -96 = 0
x(x+ 12 ) -8 (x+ 12 )= 0
(x+ 12 )(x -8 )= 0

x= -12 OR 8
Ignore negative.
so x= 8
Width 8 yards
Length 12 yards (width + 4 )
m.ananth@hotmail.ca